OpenStudy (anonymous):

A steel rod is 3 feet long. You cut three pieces that are 2 1/4 inches long and there is 1/16 waste with each cut. How much if the original rod is left?

OpenStudy (ghazi):

2 1/4= 9/4 and 1/16 total length= 9/4+1/16+x=3 x= 3- (9/4+1/16) x=3-(2.25+0.0625) x=3-2.3125 x=0.6875 feet of rod is left

OpenStudy (anonymous):

Thanks Again! :)

OpenStudy (ghazi):

:)